OpenStudy (anonymous):

Prove sin(x+y)=sin(x)cos(y) + cos(x)sin(y), use the left side to start and use the trig identities as in the proof. How do I do this?

OpenStudy (yttrium):

From the identity sin (A+b) = sinAcosB + cosAsinB We can prove that: sin(x+y) = sin(x)cos(y)+cos(x)sin(y) Hence sin(x)cos(y)+cos(x)sin(y) = sin(x)cos(y)+cos(x)sin(y) That's it. :)

OpenStudy (anonymous):

wow now that i look at it it seems so obvious haha thank you!